# Quantum-impurity sensing of altermagnetic order

### v2

@@ I Magnetostatic Green’s function and the geometric factor of the relaxation rate
In this section we present explicitly the relation between the stray magnetic field $\hat{\vec{B}}_{\rm{QI}}$ in Eq.(1) and the spin density $\hat{\vec{s}}$ and obtain explicitly the geometric factor $\mathcal{C}_{\{\Phi\}}(d,\vec{k})$ of Eq. (2). The stray magnetic field $\hat{\vec{B}}_{\rm{QI}}$ is given by the the stray magnetic field $\hat{\vec{B}}$ written in the quantum impurity (QI) principal axis frame. We keep to the setup of Fig. 1, and consider that the QI principal axis lies in the yz plane, such that $\hat{\vec{B}}_{\rm{QI}}=\mathcal{R}_{x}(\theta)\cdot\hat{\vec{B}}$, where $\mathcal{R}_{x}(\theta)$ is a rotation around the $x$ axis.
@@
$\mathcal{D}_{zz}(\vec{r},\vec{r}^{\prime})$ $=-\frac{1}{2\pi}\int d^{2}\vec{k}ke^{-k|y-y^{\prime}|}e^{i\vec{k}\cdot(\vec{\rho}-\vec{\rho}^{\prime})},$
@@ II Magnon diffusion tensor
We present here the full formula for the magnon diffusion tensor $\overleftrightarrow{D}$. We follow the procedured used for ferro- and antiferromagnets [44, 45]. Our starting point is the excess density $\delta n_{\xi}(\vec{r})$ of each magnon species $\xi=\alpha,\beta$ is
@@ II.2 Diffusion tensor
With the magnon bands we can evaluate the magnon velocities, and then calculate the magnon diffusion tensor in Eq. (S17). Even though the magnon velocities have convoluted forms, we can infer dome underlying symmetry properties that are a consequence of the symmetries of the Heisenberg Hamiltonian of Eq. (4). Those a better analyzed in a coordinate system rotated by $\pi/4$, such that
@@ IV Relaxation rate
We first recall that the geometric factor is given by $\mathcal{C}_{\{\Phi\}}(d,\vec{k})=(2\pi)^{2}k^{2}e^{-2kd}\mathcal{F}_{\{\Phi\}}(\phi_{k})$, where the function $\mathcal{F}_{\{\Phi\}}(\phi_{k})$ depends only on the orientation angles, and we work in the high-temperature regime, such that:

### v3

@@ I Magnetostatic Green’s function and the geometric factor of the relaxation rate
In this section we present explicitly the relation between the stray magnetic field $\hat{\vec{B}}_{\rm{QI}}$ in Eq.(1) and the spin density $\hat{\vec{s}}$ and obtain explicitly the geometric factor $\mathcal{C}_{\{\Phi\}}(d,\vec{k})$ of Eq. (2). Here, we indicate the set of orientation angles by $\{\Phi\}=(\theta,\varphi)$, where $\theta$ and $\varphi$ defined in the main text and in Fig. 1. The stray magnetic field $\hat{\vec{B}}_{\rm{QI}}$ is given by the the stray magnetic field $\hat{\vec{B}}$ written in the quantum impurity (QI) principal axis frame. We keep to the setup of Fig. 1, and consider that the QI principal axis lies in the yz plane, such that $\hat{\vec{B}}_{\rm{QI}}=\mathcal{R}_{x}(\theta)\cdot\hat{\vec{B}}$, where $\mathcal{R}_{x}(\theta)$ is a rotation around the $x$ axis.
@@
$\mathcal{D}_{yy}(\vec{r},\vec{r}^{\prime})$ $=-\frac{1}{2\pi}\int d^{2}\vec{k}ke^{-k|y-y^{\prime}|}e^{i\vec{k}\cdot(\vec{\rho}-\vec{\rho}^{\prime})},$
@@
To connect our results with the literature, we consider now the case in which the Fourier transform of the spin correlator $C_{\parallel}(\vec{k},\omega)$, appearing in Eq. (S4), is independent of the direction of the $k$-vector, i.e. $C_{\parallel}(\vec{k},\omega)=C_{\parallel}(k,\omega)$, where $k=|\vec{k}|$. In this case, the relaxation rate reads
$\Gamma[\omega]$ $=\gamma^{2}\tilde{\gamma}^{2}\int\frac{d^{2}\vec{k}}{(2\pi)^{2}}\mathcal{C}_{\{\Phi\}}(d,\vec{k})C_{\parallel}(k,\omega)=(\gamma\tilde{\gamma})^{2}\int\frac{dkd\varphi_{k}}{(2\pi)^{2}}k\,\mathcal{C}_{\{\Phi\}}(d,\vec{k})C_{\parallel}(k,\omega).$ (S6)
Using now Eq. (S5) we have
$\Gamma[\omega]$ $=(\gamma\tilde{\gamma})^{2}\int dkd\phi_{k}k^{3}e^{-2kd}\mathcal{F}_{\{\Phi\}}(\phi_{k})C_{\parallel}(k,\omega)$ (S7)
$=(\gamma\tilde{\gamma})^{2}\left[\int d\phi_{k}\mathcal{F}_{\{\Phi\}}(\phi_{k})\right]\left[\int dkk^{3}e^{-2kd}C_{\parallel}(k,\omega)\right]$
$=f(\{\Phi\})\left[\int dkk^{3}e^{-2kd}C_{\parallel}(k,\omega)\right],$
where
$f(\{\Phi\})=(\gamma\tilde{\gamma})^{2}\left[\int d\phi_{k}\mathcal{F}_{\{\Phi\}}(\phi_{k})\right].$ (S8)
Although being a rather long calculation, the angular integral involves only trigonometric relations. Performing it gives us
$f(\{\Phi\})=\frac{(\gamma\tilde{\gamma})^{2}\pi}{4}(5-\cos^{2}\theta+2\cos^{2}\varphi-2\cos^{2}\varphi\cos^{2}\theta).$ (S9)
This is exactly the same geometric factor shown in the SI of Ref. [36] (Eq. S16, after simple algebraic manipulations) for $\vartheta=0$, which corresponds to in-plane Néel vector, as is the case considered here.
@@ II Magnon diffusion tensor
We present here the full formula for the magnon diffusion tensor $\overleftrightarrow{D}$. We follow the procedure used for ferro- and antiferromagnets [51, 52]. Our starting point is the excess density $\delta n_{\xi}(\vec{r})$ of each magnon species $\xi=\alpha,\beta$ is
@@ II.2 Diffusion tensor
With the magnon bands we can evaluate the magnon velocities, and then calculate the magnon diffusion tensor in Eq. (S21). Even though the magnon velocities have convoluted forms, we can infer dome underlying symmetry properties that are a consequence of the symmetries of the Heisenberg Hamiltonian of Eq. (S25). Those a better analyzed in a coordinate system rotated by $\pi/4$, such that
@@ IV Relaxation rate
We first recall that the geometric factor is given by $\mathcal{C}_{\{\Phi\}}(d,\vec{k})=(2\pi)^{2}k^{2}e^{-2kd}\mathcal{F}_{\{\Phi\}}(\phi_{k})$, where the function $\mathcal{F}_{\{\Phi\}}(\phi_{k})$ depends only on the orientation angles (see Eq. (S5)), and we work in the high-temperature regime, such that: